OpenStudy (anonymous):

what is the property of 3x+7=8? I need it asap please. the answerwoud be one of these A)subtraction property of equation B)combine like terms C)division property of equation D) distribution property

Parth (parthkohli):

?

OpenStudy (amistre64):

... and its not "d"

OpenStudy (anonymous):

then what is it?

OpenStudy (amistre64):

im going to go with Parth on this one: ?

OpenStudy (amistre64):

your question makes no sense

OpenStudy (anonymous):

im suppose to choose the correct reason for the algebraic statement. in this case it for 3x+7=8

OpenStudy (amistre64):

distributive property is: \[a(b+c)=ab+ac\] which is definantly not happening; so its at least not "d".

OpenStudy (amistre64):

... the other options are an enigma. If we were to solve for the value of x: we would subtract 7 combine like terms then divide off 3 if you have material that does with this, make sure that a few of those properties its asking for arent made up.

OpenStudy (anonymous):

this is a question from my math class and i need to find out the answer, so it aint made up